# Tannakian categories: origins and summary

### v1

@@
Briefly, a tannakian category over a field $k$ is a $k$-linear abelian category with a tensor product structure having most of the properties of the category of finite-dimensional representations of an affine group scheme over $k$ except one: there need not exist an exact tensor functor to the category of $k$-vector spaces, and when one does exist there need be a canonical one. Each of the cohomology theories takes values, not just in a category of vector spaces, but in a tannakian category. For example, crystalline cohomology takes values in a category of isocrystals. These are finite-dimensional vector spaces over an extension of $\mathbb{Q}_{p}$, but only the elements of $\mathbb{Q}_{p}$ act as endomorphisms in the category. More specifically, if $\mathbb{1}$ is the unit object of the category (the tensor product of the empty set), we have $\mathrm{End}(\mathbb{1})=\mathbb{Q}_{p}$. Grothendieck’s insight is that there should be a tannakian category $\mathsf{Mot}$ over $\mathbb{Q}{}$ such that the functors to the local tannakian categories defined by the different cohomology theories factor through it. Algebraic correspondences between smooth projective algebraic varieties should define maps between motives, whose traces lie in $\mathrm{End}(\mathbb{1})=\mathbb{Q}$ and map to the traces on the various cohomology groups, which explains why the latter lie in $\mathbb{Q}$.
@@
such that, for all $k$-schemes $T$, $(S(T),G(T),(t,s),\circ)$ is a groupoid (i.e., a small category in which the morphisms are isomorphisms). A groupoid $G$ is transitive if the morphism
@@
Let $\mathrm{Aff}_{k}$ denote the category of affine $k$-schemes. For each affine $k$-scheme $S$, we let $\textsc{Fib}(\mathsf{C})_{S}$ denote the category of fibre functors of $\mathsf{C}$ over $S$. As $S$ varies, the categories $\textsc{Fib}(\mathsf{C})_{S}$ form a stack over $\mathrm{Aff}_{k}$ for the fpqc topology, and (c) of Theorem 4 implies that $\textsc{Fib}(\mathsf{C})$ is a gerb (any two fibre functors are locally isomorphic).

### v2

@@
Briefly, a tannakian category over a field $k$ is a $k$-linear abelian category with a tensor product structure having most of the properties of the category of finite-dimensional representations of an affine group scheme over $k$ except one: there need not exist an exact tensor functor to the category of $k$-vector spaces, and when one does exist there need not be a canonical one. Each of the cohomology theories takes values, not just in a category of vector spaces, but in a tannakian category. For example, crystalline cohomology takes values in a category of isocrystals. These are finite-dimensional vector spaces over an extension of $\mathbb{Q}_{p}$, but only the elements of $\mathbb{Q}_{p}$ act as endomorphisms in the category. More specifically, if $\mathbb{1}$ is the unit object of the category (the tensor product of the empty set), we have $\mathrm{End}(\mathbb{1})=\mathbb{Q}_{p}$. Grothendieck’s insight is that there should be a tannakian category $\mathsf{Mot}$ over $\mathbb{Q}{}$ such that the functors to the local tannakian categories defined by the different cohomology theories factor through it. Algebraic correspondences between smooth projective algebraic varieties should define maps between motives, whose traces lie in $\mathrm{End}(\mathbb{1})=\mathbb{Q}$ and map to the traces on the various cohomology groups, which explains why the latter lie in $\mathbb{Q}$.
@@
such that, for all $k$-schemes $T$, $(S(T),G(T),(t,s),\circ)$ is a groupoid (i.e., a small category in which the morphisms are isomorphisms). The groupoid $G$ is transitive if the morphism
@@
Let $\mathrm{Aff}_{k}$ denote the category of affine $k$-schemes. For each affine $k$-scheme $S$, we let $\textsc{Fib}(\mathsf{C})_{S}$ denote the category of fibre functors on $\mathsf{C}$ over $S$. As $S$ varies, the categories $\textsc{Fib}(\mathsf{C})_{S}$ form a stack over $\mathrm{Aff}_{k}$ for the fpqc topology, and (c) of Theorem 4 implies that $\textsc{Fib}(\mathsf{C})$ is a gerbe (any two fibre functors are locally isomorphic).